# Sloth: scaling laws for LLM skills to predict multi-benchmark performance across families

### v4

@@ Abstract
Scaling laws for large language models (LLMs) predict model performance based on parameters like size and training data. However, differences in training configurations and data processing across model families lead to significant variations in benchmark performance, making it difficult for a single scaling law to generalize across all LLMs. On the other hand, training family-specific scaling laws requires training models of varying sizes for every family. In this work, we propose Skills Scaling Laws (SSLaws, pronounced as Sloth), a novel scaling law that leverages publicly available benchmark data and assumes LLM performance is driven by low-dimensional latent skills, such as reasoning and instruction following. These latent skills are influenced by computational resources like model size and training tokens but with varying efficiencies across model families. Sloth exploits correlations across benchmarks to provide more accurate and interpretable predictions while alleviating the need to train multiple LLMs per family. We present both theoretical results on parameter identification and empirical evaluations on 12 prominent benchmarks, from Open LLM Leaderboard v1/v2, demonstrating that Sloth predicts LLM performance efficiently and offers insights into scaling behaviors for complex downstream tasks and increased test-time compute. Our code is available on https://github.com/felipemaiapolo/sloth.
Machine Learning, ICML
@@ 1 Introduction
Large Language Model (LLM) scaling laws for benchmarks and downstream tasks efficiently predict the performance of an LLM based on its parameter count and training set size. However, variations in training configurations and data processing across different model families often lead to significant differences in benchmark performance, even for models with comparable compute budgets (Ruan et al., 2024). Consequently, a single scaling law typically fails to predict performance across all LLMs accurately (Choshen et al., 2024). In contrast, creating family-specific scaling laws requires training multiple models of increasing size, which is resource-intensive.
@@
  • Demonstrating how predicted latent skills can be used to predict (i) model performance in complex downstream tasks that involve coding and emotional intelligence and (ii) LLM behavior with scaled test-time compute.
@@ 1.1 Related work
Scaling laws for deep neural networks: In recent years, researchers have studied scaling laws from different angles. Rosenfeld et al. (2019) provides experimental scaling laws that predict model loss as a function of training set size, model width, and model depth. Likewise, Kaplan et al. (2020) establishes scaling laws that primarily measure loss (perplexity) and not accuracy on downstream tasks or benchmarks. Motivated by the presence of hard limits on the size of trainable data sets but a hypothetical unlimited ability to scale models, the authors of Muennighoff et al. (2023) establish scaling laws in constrained data settings. They find that perhaps unsurprisingly, increasing computing provides diminishing returns if data does not scale. Gadre et al. (2024) addresses the gap between the assumptions in scaling laws and how training is performed in practice; in particular, they construct scaling laws that both perform well in the over-training regime and predict performance on downstream tasks. In a similar but distinct direction, some works try not only to estimate scaling laws but also respond to the following strategic question: “Given a fixed FLOPs budget, how should one trade-off model size and the number of training tokens?” For example, Hoffmann et al. (2022) provides a partial answer, introducing the celebrated family of Chinchilla scaling laws and finding that training tokens and parameter size should roughly scale together. This contrasts with the older work of Kaplan et al. (2020) that provides a series of power laws that imply that simply increasing parameter count will provide good returns. Each of these referenced works trains models with a particular pretraining setting (e.g., architecture) at various sizes and ultimately seeks to predict test loss. Our focus is distinct, we fit scaling laws on existing benchmark data of multiple model families and predict LLM benchmark performance with minimal amount of data on the new family being predicted. Even though Allen-Zhu & Li (2024), study how LLMs can retain knowledge depending on their scale, the closest related works are Owen (2024); Ruan et al. (2024); Gadre et al. (2024); we will provide a detailed comparison with their work throughout the paper.
@@
LLMs latent skills: Given that the performance of large language models (LLMs) in different and diverse benchmarks is correlated, it makes sense to think that those models have some low-dimensional latent skills that are reflected in downstream tasks. In this direction, Ilić (2023) extracts a general intelligence factor (“g-factor”) for LLMs using the Open LLM Leaderboard (Beeching et al., 2023) and GLUE (Wang et al., 2018) using factor analysis. They also verify that this “g-factor” positively correlates with model size. In a similar direction Burnell et al. (2023) uses HELM (Liang et al., 2022) data to reveal that LLM intelligence may be constituted by three distinct, yet correlated factors. They also verify a positive correlation between model size and these latent skills, yet they do not propose a formal scaling law. In their study, the authors do not account for the training set size or model family information, leading to a poor fit of the regression model; this leaves good extrapolation as an open problem we address. In Kipnis et al. (2024), a unidimensional item response theory model is applied to each one of the 6 (filtered) benchmarks of the Open LLM Leaderboard. A factor analysis on the skill parameters shows that the main factor (carrying  80% of the data variability) is highly correlated with the “grand” (average) score of LLMs. In a related but different direction, Maia Polo et al. (2024a, b) show that inferring low-dimensional latent skills of LLMs can make model evaluation much more efficient, saving up to 140x in computing power. In this work, we explicitly model LLM skills as a function of computing resources, which enables the creation of accurate and interpretable scaling laws for benchmark performances.
@@ 2.1 Problem Statement
$\mu_{ij}(s,t)={\mathbf{E}}[Y_{ij}(s,t)].$ (2.1)
Here, ${\mathbf{E}}[\cdot]$ should be interpreted as a central tendency summary measure of a random variable, such as the mean or median. Ideally, the model for $\mu_{ij}$ will be simple and some of its parameters will be shared among model families and benchmarks; in this case, the model becomes more interpretable and more data can be used in the fitting process, making the model better estimated. From now on, we denote the set of model families as =$\mathcal{I}=\{1,\cdots,I\}$ and the set of benchmarks as $\mathcal{J}=\{1,\cdots,J\}$.
@@ 2.2 Previous approaches to scaling laws for benchmarks
The closest works to ours that propose models for $\mu_{ij}(s,t)$ (2.1) are Owen (2024); Ruan et al. (2024), and Gadre et al. (2024). While Gadre et al. (2024) indirectly model the quantity of interest via the LLMs perplexity in specific datasets, which might not be readily available, Owen (2024) and Ruan et al. (2024) model $\mu_{ij}(s,t)$ directly through a regression model connecting compute and benchmark performance. One assumption they made is that the performance on benchmarks only depends on $s$ and $t$ through the total amount of training FLOPs, which can be approximated by $c(s,t)=6st$. That is, if $\sigma:{\mathbf{R}}\to[0,1]$ denotes a fixed activation function, e.g., the standard logistic (sigmoid) function, and $\gamma_{j}\in[0,1]$, then it is assumed that
@@
where $\eta_{ij}:{\mathbf{R}}^{2}\to{\mathbf{R}}$ denotes a linear predictor such that $\eta_{ij}(s,t)=\alpha_{ij}+\beta_{ij}\log c(s,t)$, which can be easily interpreted. Here, $\gamma_{j}$ adjusts the lower asymptote of $\mu_{ij}$ and accounts for the probability of LLMs scoring correctly by chance. Owen (2024), in their best performing models, considers the case in which $\gamma_{j}=0$ (or adds a similar offset parameter to the model) and the parameters $\alpha_{ij}$ and $\beta_{ij}$ are independent of the model family $i$. On the other hand, Ruan et al. (2024) consider both $\alpha_{ij}$ and $\beta_{ij}$ to be family-dependent and, in their most general model, $\gamma_{j}$ can assume values in $[0,1]$.
The biggest issue with previous approaches when modeling $\mu_{ij}$ is that they are either too restrictive or too flexible. From the restrictive side, they assume that (i) $\mu_{ij}$ depends on $s$ and $t$ only through FLOPs, (ii) there are no family-dependent parameters, or (iii) the activation function $\sigma$ is fixed and well-specified. From the flexibility side, Ruan et al. (2024) assume both $\alpha_{ij}$ and $\beta_{ij}$ to be family dependent making estimation hard (or impossible) depending on the number of models we see for each family. From Ruan et al. (2024): “(…) fitting such a scaling law can be tricky, as each model family $f$ and downstream benchmark has its own scaling coefficients $\beta_{f}$ and $\alpha_{f}$. This means that scaling experiments, especially for post-training analysis, are often fitted on very few (3-5) models sharing the same model family (…).” Thus, in their experiments, they consider a different problem setting, where a large LLM has been trained and evaluated on some benchmarks and use their method to predict its performance on other benchmarks.
At the end of the day, Owen (2024) and Ruan et al. (2024) end up working in different setups: Owen (2024) does not use family information at prediction time, making their scaling law less accurate but more generalizable, and Ruan et al. (2024) assume families are important at prediction time but consider that the target model has already been trained, making their scaling law less applicable in practice and more interesting from an interpretability point of view. In this work we wish to instead predict the performance of a larger LLM without having to train it but taking family information into account, thus allowing practitioners to make decisions regarding investing resources into training large LLMs. Moreover, our formulation also allows interpretable insights from the data. Despite different setups, we make comparisons with Owen (2024) and Ruan et al. (2024) throughout this work by considering their applications/adaptations as baselines.
@@ 3.1 Model architecture
Inspired by the latent skills (e.g., reasoning, language modeling, instruction following) inferred from benchmark data in Burnell et al. (2023); Ilić (2023); Ruan et al. (2024); Gor et al. ; Kipnis et al. (2024); Maia Polo et al. (2024a, b), we propose creating a scaling law for LLMs skills by leveraging the correlation structure of the benchmarks; for example, we model how the construct “reasoning” scales with compute instead of modeling benchmarks scores directly. The two major advantages of this approach are better performance prediction since we have fewer parameters to fit (reducing overfitting) and extra interpretability/insights. Concretely, we model $\eta_{ij}(s,t)$’s simultaneously for benchmarks $j\in\mathcal{J}$ as each being a linear combination of the same low-dimensional latent skills $\theta_{i}(s,t)\in{\mathbf{R}}^{d}$ plus a bias term $b\in{\mathbf{R}}^{J}$, where $d\ll J=|\mathcal{J}|$. Denote $\eta_{i}(s,t)\in{\mathbf{R}}^{J}$ as the vector of $\{\eta_{ij}(s,t)\}_{j\in\mathcal{J}}$. Mathematically, we have
@@
One can see that $\Lambda\in{\mathbf{R}}^{J\times d}$ encodes the correlation structure between the benchmarks; in particular, it tells us which benchmark measures overlapping (or distinct) skills. Interestingly, our model has a strong connection with factor analysis (FA) models, which we elaborate on in detail in Appendix D. In FA, the matrix $\Lambda$ is known as factor loadings while $\theta_{i}(s,t)$ are known as factors.
Next, we propose a model for $\theta_{i}(s,t)$. Inspired by models used in Economics, we use the family of translog production functions from stochastic frontier analysis (Kumbhakar & Lovell, 2003):
@@
Note that (i) the intercept parameter $\alpha_{ik}$ is indeed family-dependent while each skill slope is shared across families and (ii) $\theta_{i}$ can depend on $s$ and $t$ not only through $c(s,t)$. In economic terms, the intercept term $\alpha_{ik}$ can be interpreted as an efficiency measure of the family $i$ in converting compute to performance for skill $k$ and, in practice, will absorb all hidden factors specific to family $i$ such as data quality, post-training factors, etc.. We note that the interaction term in ($\log(s)\log(t)$) accounts for the fact that the impact of $\log(s)$ and $\log(t)$ on skills might depend on each other; in Appendix E, we show some evidence that this is indeed the case. Additionally to the changes in $\eta_{ij}$, we propose making the activation function $\sigma$ trainable and specific to each benchmark $j$ if needed. To that end, we adopt a semi-parametric single-index model using neural networks (Bietti et al., 2022). To make the results more behaved if (out-of-support) generalization is needed, we assume $\sigma_{j}:{\mathbf{R}}\to[0,1]$ is given by a monotonic (increasing) neural network, which can be achieved by constraining its weights to be non-negative (Sill, 1997) and its last activation function to be sigmoid. We note, however, that one can always forgo training of the link function and instead assume a sigmoid structure as this simplifies model fitting and may make the model more interpretable. We give more details about model fitting in Appendix C.
@@ 3.2 Interpretability and practical considerations post model fitting
Ideally, to make reasonable interpretations of models like Sloth, we need its parameters to be identifiable, i.e., the map connecting $\eta_{ij}(s,t)$’s and model parameters should be bijective. Unfortunately, as in all exploratory factor models, this is not the case. However, we theoretically show in Appendix B that the model parameters are identifiable up to some parameter transformations. This observation allows us to find a configuration that makes the model more interpretable by mirroring a standard approach used in factor analysis, e.g., in Chen et al. (2019)’s applications. First, we fit Sloth without any constraints on the covariance of skills obtaining the estimates $(\hat{\Lambda},\hat{b},\hat{B})$. Second, we find the matrix $A\in{\mathbf{R}}^{d\times d}$ such that the skills $X\hat{B}A$ have covariance identity, update $\hat{B}\leftarrow\hat{B}A$, and update $\hat{\Lambda}\leftarrow\hat{\Lambda}(A^{\top})^{-1}$ so the model outputs remains unchanged, because $\hat{\Lambda}(X\hat{B})^{\top}=\hat{\Lambda}(A^{\top})^{-1}(X\hat{B}A)^{\top}$. Third, we find a matrix $M\in{\mathbf{R}}^{d\times d}$ such that $\hat{\Lambda}M$ is easily interpretable (e.g., it is a sparse matrix); there are different methods to find $M$ and, in this paper, we use the Geomin (Yates, 1987; Chen et al., 2019) oblique rotation method to find a suitable $M$ using the Python package FactorAnalyzer (Biggs, 2019). We then update $\hat{\Lambda}\leftarrow\hat{\Lambda}M$ and, to make the model invariant, we also update $\hat{B}\leftarrow\hat{B}(M^{\top})^{-1}$; the resulting skills are still guaranteed to have unitary standard deviations, so their covariance equals their correlation. Finally, we standardize the columns of the skills $X\hat{B}$ to have zero mean, while keeping the correlation structure unchanged. This last step implies that $\hat{b}$ must be translated to make the model invariant.

### v5

@@ Abstract
Scaling laws for large language models (LLMs) predict model performance based on parameters like size and training data. However, differences in training configurations and data processing across model families lead to significant variations in benchmark performance, making it difficult for a single scaling law to generalize across all LLMs. On the other hand, training family-specific scaling laws requires training models of varying sizes for every family. In this work, we propose Skills Scaling Laws (SSLaws, pronounced as Sloth), a novel scaling law that leverages publicly available benchmark data and assumes LLM performance is driven by low-dimensional latent skills, such as reasoning and instruction following. These latent skills are influenced by computational resources like model size and training tokens, but with varying efficiencies across model families. Sloth exploits correlations across benchmarks to provide more accurate and interpretable predictions while alleviating the need to train multiple LLMs per family. We present both theoretical results on parameter identification and empirical evaluations on 12 prominent benchmarks, from Open LLM Leaderboard v1/v2, demonstrating that Sloth predicts LLM performance accurately and offers insights into scaling behaviors for complex downstream tasks, increased test-time compute, and compute-optimal scaling of skills. Our code can be found on https://github.com/felipemaiapolo/sloth.
@@ 1 Introduction
Large Language Model (LLM) scaling laws for benchmarks and downstream tasks efficiently predict the performance of an LLM based on its parameter count and training set size. However, variations in training configurations and data processing across different model families often lead to significant differences in benchmark performance, even for models with comparable compute budgets (ruan2024observational). Consequently, a single scaling law typically fails to predict performance across all LLMs accurately (choshen2024hitchhiker). In contrast, creating family-specific scaling laws requires training multiple models of increasing size, which is resource-intensive.
@@
  • Demonstrating how predicted latent skills and our model can be used to (i) predict model performance in complex downstream tasks that involve coding and emotional intelligence, (ii) predict LLM behavior with scaled test-time compute, and (iii) derive optimal-scaling rules for skills.
@@ 1.1 Related work
Scaling laws for deep neural networks: In recent years, researchers have studied scaling laws from different angles. rosenfeld2019constructivepredictiongeneralizationerror provides experimental scaling laws that predict model loss as a function of training set size, model width, and model depth. Likewise, kaplan2020scalinglawsneurallanguage establishes scaling laws that primarily measure loss (perplexity) and not accuracy on downstream tasks or benchmarks. Motivated by the presence of hard limits on the size of trainable data sets but a hypothetical unlimited ability to scale models, the authors of muennighoff2023scalingdataconstrainedlanguagemodels establish scaling laws in constrained data settings. They find that perhaps unsurprisingly, increasing computing provides diminishing returns if data does not scale. gadre2024languagemodelsscalereliably addresses the gap between the assumptions in scaling laws and how training is performed in practice; in particular, they construct scaling laws that both perform well in the over-training regime and predict performance on downstream tasks. In a similar but distinct direction, some works try not only to estimate scaling laws but also respond to the following strategic question: “Given a fixed FLOPs budget, how should one trade-off model size and the number of training tokens?” For example, hoffmann2022training provides a partial answer, introducing the celebrated family of Chinchilla scaling laws and finding that training tokens and parameter size should roughly scale together. This contrasts with the older work of kaplan2020scalinglawsneurallanguage that provides a series of power laws that imply that simply increasing parameter count will provide good returns. Each of these referenced works trains models with a particular pretraining setting (e.g., architecture) at various sizes and ultimately seeks to predict test loss. Our focus is distinct, we fit scaling laws on existing benchmark data of multiple model families and predict LLM benchmark performance with minimal amount of data on the new family being predicted. Even though allen2024physics, study how LLMs can retain knowledge depending on their scale, the closest related works are owen2024predictable; ruan2024observational; gadre2024languagemodelsscalereliably; we will provide a detailed comparison with their work throughout the paper.
@@
LLMs latent skills: Given that the performance of large language models (LLMs) in different and diverse benchmarks is correlated, it makes sense to think that those models have some low-dimensional latent skills that are reflected in downstream tasks. In this direction, ilic2023unveiling extracts a general intelligence factor (“g-factor”) for LLMs using the Open LLM Leaderboard (openllmldrboard) and GLUE (wang2018glue) using factor analysis. They also verify that this “g-factor” positively correlates with model size. In a similar direction burnell2023revealing uses HELM (liang2022holistic) data to reveal that LLM intelligence may be constituted by three distinct, yet correlated factors. They also verify a positive correlation between model size and these latent skills, yet they do not propose a formal scaling law. In their study, the authors do not account for the training set size or model family information, leading to a poor fit of the regression model; this leaves good extrapolation as an open problem we address. In kipnis2024texttt, a unidimensional item response theory model is applied to each one of the 6 (filtered) benchmarks of the Open LLM Leaderboard. A factor analysis on the skill parameters shows that the main factor (carrying  80% of the data variability) is highly correlated with the “grand” (average) score of LLMs. In a related but different direction, polo2024tinybenchmarks; polo2024efficient show that inferring low-dimensional latent skills of LLMs can make model evaluation much more efficient, saving up to 140x in computing power. In this work, we explicitly model LLM skills as a function of computing resources, which enables the creation of accurate and interpretable scaling laws for benchmark performances.
@@ 2.1 Problem Statement
$\mu_{ij}(s,t)=m[Y_{ij}(s,t)].$ (2.1)
Here, $m[\cdot]$ should be interpreted as a central tendency summary measure of a random variable, such as the mean or median [^1] Because we minimize the Huber loss in this paper, we aim to approximate the median. Ideally, the model for $\mu_{ij}$ will be simple and some of its parameters will be shared among model families and benchmarks; in this case, the model becomes more interpretable and more data can be used in the fitting process, making the model better estimated. From now on, we denote the set of model families as $\mathcal{I}=\{1,\cdots,I\}$ and the set of benchmarks as $\mathcal{J}=\{1,\cdots,J\}$.
@@ 2.2 Previous approaches to scaling laws for benchmarks
The closest works to ours that propose models for $\mu_{ij}(s,t)$ (2.1) are owen2024predictable; ruan2024observational, and gadre2024languagemodelsscalereliably. While gadre2024languagemodelsscalereliably indirectly model the quantity of interest via the LLMs perplexity in specific datasets, which might not be readily available, owen2024predictable and ruan2024observational model $\mu_{ij}(s,t)$ directly through a regression model connecting compute and benchmark performance. One assumption they made is that the performance on benchmarks only depends on $s$ and $t$ through the total amount of training FLOPs, which can be approximated by $c(s,t)=6st$. That is, if $\sigma:{\mathbf{R}}\to[0,1]$ denotes a fixed activation function, e.g., the standard logistic (sigmoid) function, and $\gamma_{j}\in[0,1]$, then it is assumed that
@@
where $\eta_{ij}:{\mathbf{R}}^{2}\to{\mathbf{R}}$ denotes a linear predictor such that $\eta_{ij}(s,t)=\alpha_{ij}+\beta_{ij}\log c(s,t)$, which can be easily interpreted. Here, $\gamma_{j}$ adjusts the lower asymptote of $\mu_{ij}$ and accounts for the probability of LLMs scoring correctly by chance. owen2024predictable, in their best performing models, considers the case in which $\gamma_{j}=0$ (or adds a similar offset parameter to the model) and the parameters $\alpha_{ij}$ and $\beta_{ij}$ are independent of the model family $i$. On the other hand, ruan2024observational consider both $\alpha_{ij}$ and $\beta_{ij}$ to be family-dependent and, in their most general model, $\gamma_{j}$ can assume values in $[0,1]$.
The biggest issue with previous approaches when modeling $\mu_{ij}$ is that they are either too restrictive or too flexible. From the restrictive side, they assume that (i) $\mu_{ij}$ depends on $s$ and $t$ only through FLOPs, (ii) there are no family-dependent parameters, or (iii) the activation function $\sigma$ is fixed and well-specified. From the flexibility side, ruan2024observational assume both $\alpha_{ij}$ and $\beta_{ij}$ to be family dependent making estimation hard (or impossible) depending on the number of models we see for each family. From ruan2024observational: “(…) fitting such a scaling law can be tricky, as each model family $f$ and downstream benchmark has its own scaling coefficients $\beta_{f}$ and $\alpha_{f}$. This means that scaling experiments, especially for post-training analysis, are often fitted on very few (3-5) models sharing the same model family (…).” Thus, in their experiments, they consider a different problem setting, where a large LLM has been trained and evaluated on some benchmarks and use their method to predict its performance on other benchmarks.
At the end of the day, owen2024predictable and ruan2024observational end up working in different setups: owen2024predictable does not use family information at prediction time, making their scaling law less accurate but more generalizable, and ruan2024observational assume families are important at prediction time but consider that the target model has already been trained, making their scaling law less applicable in practice and more interesting from an interpretability point of view. In this work we wish to instead predict the performance of a larger LLM without having to train it but taking family information into account, thus allowing practitioners to make decisions regarding investing resources into training large LLMs. Moreover, our formulation also allows interpretable insights from the data. Despite different setups, we make comparisons with owen2024predictable and ruan2024observational throughout this work by considering their applications/adaptations as baselines.
@@ 3.1 Model architecture
Inspired by the latent skills (e.g., reasoning, language modeling, instruction following) inferred from benchmark data in burnell2023revealing; ilic2023unveiling; ruan2024observational; gorgreat; kipnis2024texttt; polo2024tinybenchmarks; polo2024efficient, we propose creating a scaling law for LLMs skills by leveraging the correlation structure of the benchmarks; for example, we model how the construct “reasoning” scales with compute instead of modeling benchmarks scores directly. The two major advantages of this approach are better performance prediction since we have fewer parameters to fit (reducing overfitting) and extra interpretability/insights. Concretely, we model $\eta_{ij}(s,t)$’s simultaneously for benchmarks $j\in\mathcal{J}$ as each being a linear combination of the same low-dimensional latent skills $\theta_{i}(s,t)\in{\mathbf{R}}^{d}$ plus a bias term $b\in{\mathbf{R}}^{J}$, where $d\ll J=|\mathcal{J}|$. Denote $\eta_{i}(s,t)\in{\mathbf{R}}^{J}$ as the vector of $\{\eta_{ij}(s,t)\}_{j\in\mathcal{J}}$. Mathematically, we have
@@
One can see that $\Lambda\in{\mathbf{R}}^{J\times d}$ encodes the correlation structure between the benchmarks; in particular, it tells us which benchmark measures overlapping (or distinct) skills. Interestingly, our model has a strong connection with factor analysis (FA) models, which we elaborate on in detail in Appendix C. In FA, the matrix $\Lambda$ is known as factor loadings while $\theta_{i}(s,t)$ are known as factors.
Next, we propose a model for $\theta_{i}(s,t)$. Inspired by models used in Economics, we use the family of translog production functions from stochastic frontier analysis (kumbhakar2003stochastic):
@@
Note that (i) the intercept parameter $\alpha_{ik}$ is indeed family-dependent while each skill slope is shared across families and (ii) $\theta_{i}$ can depend on $s$ and $t$ not only through $c(s,t)$. In economic terms, the intercept term $\alpha_{ik}$ can be interpreted as an efficiency measure of the family $i$ in converting compute to performance for skill $k$ and, in practice, will absorb all hidden factors specific to family $i$ such as data quality, post-training factors, etc.. We note that the interaction term in ($\log(s)\log(t)$) accounts for the fact that the impact of $\log(s)$ and $\log(t)$ on skills might depend on each other; in Appendix D, we show some evidence that this is indeed the case. Additionally to the changes in $\eta_{ij}$, we propose making the activation function $\sigma$ trainable and specific to each benchmark $j$ if needed. To that end, we adopt a semi-parametric single-index model using neural networks (bietti2022learning). To make the results more behaved if (out-of-support) generalization is needed, we assume $\sigma_{j}:{\mathbf{R}}\to[0,1]$ is given by a monotonic (increasing) neural network, which can be achieved by constraining its weights to be non-negative (sill1997monotonic) and its last activation function to be sigmoid. We note, however, that one can always forgo training of the link function and instead assume a sigmoid structure as this simplifies model fitting and may make the model more interpretable. We give more details about model fitting in Section 3.2. Since Sloth is a simple neural network, both model fitting and prediction are done within seconds by a commercial laptop.
3.2 Model fitting
[…]
where $\ell_{\delta}$ is given by the Huber loss with hyperparameter $\delta=.01$ and $\mu_{ij}(s,t)$ denotes the most general version of our model. We minimize the loss function via gradient descent using the Adam optimizer (kingma2017Adam) with a decaying learning rate. We parameterize $\gamma_{j}$ using the sigmoid transformation to guarantee the constraints are satisfied. Similarly, we truncate the weights of the two-hidden-layer neural network $\sigma_{j}$ to ensure the trainable function is increasing. If one desires, $\sigma_{j}$’s can be set to fixed functions, e.g., sigmoid, and $\gamma_{j}$’s can be fixed beforehand. Unfortunately, the minimization problem is not convex as expected when fitting factor-analysis-like models; multiple initializations of the optimizer can be applied to guarantee a better fit.
@@ 3.3 Interpretability and practical considerations post model fitting
Ideally, to make reasonable interpretations of models like Sloth, we need its parameters to be identifiable, i.e., the map connecting $\eta_{ij}(s,t)$’s and model parameters should be bijective. Unfortunately, as in all exploratory factor models, this is not the case. However, we theoretically show in Appendix A that the model parameters are identifiable up to some parameter transformations. This observation allows us to find a configuration that makes the model more interpretable by mirroring a standard approach used in factor analysis, e.g., in chen2019joint’s applications. The main idea is that we fit Sloth without any constraints and then find a configuration of skills (using factor rotation) that makes results interpretable. We detail the applied process in Appendix A.1.